In a constant‑pressure calorimeter, 60.0 mL of 0.330 M Ba(OH)2 was added to 60.0 mL of 0.660 M HCl.
The reaction caused the temperature of the solution to rise from 23.65 ∘C to 28.15 ∘C. If the solution has the same density and specific heat as water ( 1.00 g/mL and 4.184J/g⋅°C,) respectively), what is Δ
1
Expert's answer
2020-05-19T08:42:35-0400

H+ + OH- = H2O

n(H+) = n(OH-) = 0.660 ∙ 60 / 1000 = 0.0396

ΔrH = 4.184 ∙ 120 ∙ (28.15 - 23.65) / 0.0396 = 57054 J ∙ mol-1 ∙ K-1 = 57.05 kJ ∙ mol-1 ∙ K-1

Answer: ΔrH = 57.05 kJ ∙ mol-1 ∙ K-1.